# The Wisdom of a Crowd of Brains: A Universal Brain Encoder

### v2

@@ 1 Introduction
fMRI (functional MRI) has emerged as a powerful tool for measuring brain activity. This enables brain scientists to explore active brain areas during various functions and behaviors (Kanwisher et al., 1997; Epstein & Kanwisher, 1998; Downing et al., 2001; Tang et al., 2017; Heeger & Ress, 2002). However, a human can spend only limited time inside an fMRI machine. This results in fMRI-datasets too small to span the huge space of brain functionality or visual stimuli (natural images). Moreover, the variability in brain structure and function responses between different people (Riddle & Purves, 1995; Frost & Goebel, 2012; Conroy et al., 2013; Zhen et al., 2015) makes it difficult to combine data across individuals that have not been exposed to the same stimuli. All of these form severe limitations on current ability to analyze brain functionality.
@@
Image-to-fMRI encoding models, which predict fMRI responses to natural images, have greatly advanced the field. With the rise of deep learning, sophisticated encoding models have emerged  (Yamins et al., 2014; Eickenberg et al., 2017; Wen et al., 2018c, a; Beliy et al., 2019; Gaziv et al., 2022), offering novel insights into brain function (Tang et al., 2024; Henderson et al., 2023; Gu et al., 2023; Luo et al., 2024, 2023). However, these models are primarily subject-specific and machine-specific, requiring extensive individual data (which is prohibitive) for effective training. This limits the practical use of existing brain-encoders, and prevents their ability to leverage cross-subject data. Attempts to create multi-subject encoders (e.g., (Van Uden et al., 2018; Khosla et al., 2020; Wen et al., 2018b; Gu et al., 2022)) have so far been very restrictive (see Sec. 2), showing no generalization across different datasets/machines.
@@ 2 Related Work
Methods for Mapping visual stimuli to brain activity (“Image-to-fMRI Encoding”), have significantly advanced over the years, with contributions to Neuroscience. Initially, these models utilized linear regression between hand-crafted image features, to predict fMRI responses on images (Kay et al., 2008; Naselaris et al., 2011). Over time, the field has evolved to incorporate deep learning approaches both for image feature extraction and training (Yamins et al., 2014; Eickenberg et al., 2017; Wen et al., 2018c, a; Beliy et al., 2019; Gaziv et al., 2022; Wang et al., 2023; Takagi & Nishimoto, 2023). These models are typically subject-specific, due to substantial differences between brain responses of different people (Riddle & Purves, 1995; Frost & Goebel, 2012; Conroy et al., 2013; Zhen et al., 2015). This restricts the encoder to subject-specific data, and limits its generalization across subjects.
Multi-Subject Brain Models: Efficiently integrating data from multiple subjects is challenging due to anatomical and functional differences between brains. While anatomical alignment (Mazziotta et al., 2001; Talairach, 1988; Fischl, 2012; Dale et al., 1999) aligns brains to a common anatomical space, it fails to provide accurate functional correspondences(Mazziotta et al., 2001; Haxby et al., 2011; Yamada et al., 2015; Brett et al., 2002; Wasserman et al., 2024). Although we deal here with Multi-Subject Brain Encoding, we review both Multi-Subject Encoding & Decoding.
@@
In fMRI-to-Image decoding, the goal is to reconstruct an image given its corresponding fMRI brain recording. Recent advanced fMRI decoders leverage multi-subject data (e.g., ‘MindBridge’ (Wang et al., 2024), ‘MindEye2’ (Scotti et al., 2024b), (Gong et al., 2024) and (Han et al., 2024)), by projecting the subject-specific fMRIs of all subjects into a shared space, which is then fed into a shared image-decoding network. However, the problems of multi-subject fMRI Decoding & Encoding are fundamentally different. In multi-subject Decoding, the network inputs reside in different spaces (individual subject-specific fMRIs), whereas the network outputs reside in a shared space (the Image-Space). Thus projecting to a shared space before decoding is a valid solution. In contrast, in multi-subject Encoding, the inputs (images) reside in a shared space, whereas the outputs (individual subject-specific fMRIs) reside in completely different spaces. Thus the above approach of ‘MindBridge’&‘MindEye2’ will not apply to multi-subject Encoding, as projecting to a shared space will lose most subject-specific fMRI variability (output variability). This is resolved by our new voxel-centric model.
@@
Few attempts have been made to leverage multi-subject data for Image-to-fMRI Encoding. Van Uden et al. (2018) integrated data from multiple subjects, but required them to be exposed to same images, which is very restrictive. Our approach requires NO shared data between subjects. Wen et al. (2018b) and Gu et al. (2022) finetune models of one subject to another. This allows per subject adaption, but largely limits the information sharing as they are not trained jointly. Lastly, Khosla et al. (2020) use a model with part shared and part subject-specific. However, to obtain good results they predict only a small subset of voxels which have high correlation across brains.
@@ 3.1 Overview of the Approach
The shared network components (shared across all brain-voxels of all subjects) include the feature extraction block (Fig2a), and the voxel-image cross-attention block (Fig.2c). Given any image, our encoder extracts from it various image features (from low-level to high-level) using a DINO-v2 (Oquab et al., 2023) adapted model as our feature extraction block. It outputs image features from different intermediate layers of DINO, allowing each brain-voxel to attend to the appropriate semantic levels of features that align with its functionality. These image features, along with a specific voxel-embedding, are processed through the cross-attention block, to effectively integrate and predict the voxel’s activation in response to the given image. Note that this architecture is indifferent to the number of brain-voxels in each fMRI scan, hence applies to any fMRI scan.
@@ 3.2 Architecture and Training
(Fig. 2a). This block utilizes an adapted DINO model (Oquab et al., 2023) to derive multi-scale image features. Features are extracted from L=5 intermediate layers of the DINO-v2 VIT-L/14 model (layers 1,6,12,18,24), where lower layers capture low-level image features and higher layers provide more semantic information. This hierarchical feature extraction is crucial, as voxels in the visual cortex correspond to a range of image attributes – from simple visual details to complex semantic content. Each layer’s features are projected to a lower-dimension C (via a linear layer), and are then concatenated along another dimension of length L (the number of layers). Since DINO operates on P image patches, the final feature output is of size L×P×C. In order to transform Dino features into features suitable for predicting brain activity, we used a LoRA inspired approach (Hu et al., 2021), that is more suitable for data-limited settings (see A.2 for details).
@@
(iii) Lastly, the Functional-attention performs a weighted summation of the spatially-attended features to derive a single scalar voxel activation. In this layer, $\mathbb{v}$ represents the flattened MLP output (size 1x(L*C) ), $\mathbb{q}$ is the voxel embedding itself, and K is learned functional-embedding that has an entry for each of the LXC features (size (L*C)$\times$×E). The output is calculated via $(\mathbb{q}K^{T})\mathbb{v}^{T}$. This block outputs the voxel prediction as a scalar value.
@@ 4 Experiments & Evaluations
We experimented with 3 prominent fMRI datasets: (i) The old “vim1” dataset (Kay et al., 2008; Naselaris et al., 2009; Kay et al., 2011), which has 1750 train and 120 test grayscale images, and their corresponding 4-Tesla fMRI scans for 2 subjects. (ii) The “fMRI-on-ImageNet” scans (Horikawa & Kamitani, 2017b) comprising 1200 train and 50 test pairs of natural images from ImageNet, with 3-Tesla fMRI recordings on 5 subjects. (iii) The ”Natural Scenes Dataset” (NSD) (Allen et al., 2022), a new 7-Tesla dataset with 8 subjects, each having around 9000 unique subject-specific images, and $\sim$1000 images shared across all subjects (which we use as our test set). This results in a total of 73,000 images, all taken from the COCO dataset (Deng et al., 2009). A few example images from each of the 3 datasets are displayed in Fig. 1. A subset of visually sensitive brain voxels was either provided by the dataset or selected based on SNR (see Appendix A.1)
@@
Figure 4: The Wisdom of a Crowd of Brains. By aggregating data from multiple subjects, our Universal-Encoder improves encoding of any individual subject. We compared 3 models: (i) ”Baseline” single-subject encoder of (Gaziv et al., 2022), (ii) ”Universal Encoder - single subject” – our architecture trained on each subject separately, (iii) ”Universal Encoder - multiple subjects” – our model trained on data from 8 subjects. (a) Pearson Correlation (per voxel) between predicted & ground-truth fMRI (Median value, 75th & 25th percentiles). (b) Retrieval Accuracy (Top-1 & Top-5) of the GT image per “Query” fMRI.
@@ 4.1 The Wisdom of  a “Crowd of Brains”
We first demonstrate the Universal-Encoder’s ability to exploit data from multiple subjects, without any shared-data. For this we use the 7-Tesla NSD dataset. We train on union of all 8 subject-specific images ($\sim$72,000 non-shared data), and test on the $\sim$1,000 shared images (the images that all 8 subjects saw). We compared 3 models in our evaluation: (i) As a baseline, we used the image-to-fMRI encoder of (Gaziv et al., 2022), trained separately for each subject on their subject-specific training-set (“Baseline encoder”). (ii) Our Universal-Encoder trained on each subject separately (“Universal Encoder - single subject”), and (iii) Our Universal-Encoder trained on all 8 subjects jointly, using their combined training-sets, and tested on each subject individually (“Universal Encoder - multiple subjects”). A comparison to two other prominent encoder models (Wang et al., 2023; Takagi & Nishimoto, 2023) (which have lower performance), is further provided in the appendix Fig. S8.
Fig. 4a shows the median Pearson correlation value (along with the $25$th & $75$th percentiles), computed between the predicted activations and the ground-truth activations for all fMRI voxels. These are computed per subject, for each of the 3 models. Our Universal-Encoder, even when trained on a single subject, performs consistently better than the Baseline-encoder. Moreover, when our Universal-Encoder is trained jointly on the training sets of all subjects, it consistently outperforms all subject-specific models. It obtains notable improvements for both the “best” subjects (Subject1: $\sim$7improvement) and the “worst” subject (Subject8: $\sim$15improvement). Statistical significance of the performance gap is assessed in Appendix C, confirming significant superiority of the Universal-Encoder. Per-region results are further presented in Fig. S10.
@@
Fig. 4b presents quantitative Retrieval results evaluated per subject for all 3 models. It shows the percent of times the correct image (corresponding to the ”Query” fMRI) was ranked $1$st (Top-1) and among the Top-5 retrieved images. The results indicate superior performance of the multi-subject Universal-Encoder compared to the 2 other models in both Top-1 & Top-5. Statistical significance of the Universal Encoder’s improvement over other models is verified in Appendix C & Table.S2. Our experiments demonstrate the model’s ability to effectively aggregate data from multiple subjects (who viewed different images), while enhancing the performance of each subject individually.
@@ 4.2 The Wisdom of  a “Crowd of Datasets”
Fig. 5a compares 3 models for the “fMRI-on-ImageNet” dataset: (i) Single-subject encoder of (Gaziv et al., 2022) (“Baseline encoder”), (ii) Our Universal-Encoder trained on all subjects in “fMRI-on-ImageNet” (“Universal Encoder - same dataset”), and (iii) Universal-Encoder trained on subjects from both “fMRI-on-ImageNet” and NSD (“Universal Encoder - multi datasets”). Our multi-subject same-dataset encoder (green) outperforms the single-subject baseline model (blue). Adding data from NSD yields even further improvement (red). Median correlation, 75th & 25th percentiles are shown. Fig. 5b shows results for VIM1 dataset. Here too, adding training data from NSD significantly enhances encoding performance on VIM1.
@@ 4.3 Transfer-Learning to New Subjects/Datasets
The transferred Universal-Encoder significantly outperforms any single-subject model on all datasets. Transfer-Learning to Subject1 within NSD (Fig. 6a) obtains $\geq$77improvement for any number of training examples. Moreover, as few as 100 subject-specific training examples suffice to obtain better results than a subject-specific model trained on the entire train-set (9000 examples). A similar gap in performance is achieved for Subject8.
@@
Fig. 6b (”fMRI-on-ImageNet”) & Fig. 6c (VIM1) demonstrate transfer learning from a new 7-Tesla dataset to older lower-resolution 3T or 4T datasets. These were scanned on very different types of images (e.g., VIM1 was scanned on B/W images with a circular black mask), and have much smaller train-sets. Transfer-Learning shows a significant improvement for any number of training data, with an improvement $\geq$84for ‘fMRI-on-ImageNet’ & $\sim$45for VIM1, using the entire (small) subject-specific train-set (significance testing over model initialization in Table. S3).
@@ Appendix A Further Technical Details
The datasets utilized in our study comprise BOLD fMRI responses to various natural images, recorded over multiple scanning sessions. Peak BOLD responses corresponding to each stimulus were estimated, resulting in a scalar value for each voxel for each image. Each dataset underwent unique pre-processing procedures, detailed in their respective publications (Kay et al., 2008; Naselaris et al., 2009; Horikawa & Kamitani, 2017a; Allen et al., 2022).
@@
For the NSD dataset (Allen et al., 2022), we used a post-processed dataset with voxel selection from (Gifford et al., 2023). They chose vision-related areas, resulting in a total of approximately 40,000 voxels, and provided voxels after per voxel normalization.
@@
For the fMRI-on-ImageNet dataset (Horikawa & Kamitani, 2017b), a relevant set of around 5,000 voxels was already provided. We implemented Z-scoring normalization ourselves for this dataset.
@@
For the VIM1 dataset (Kay et al., 2008; Naselaris et al., 2009), we selected the best 7,000 voxels according to the highest Signal-to-Noise Ratio (SNR). SNR is calculated as the ratio of the variance of averaged (repeated) measurements for different stimuli to the average variance of measurements for the same stimuli. This approach ensures the selection of voxels most representative of neural activity in response to diverse visual stimuli. Z-scoring normalization was also implemented for this dataset.

### v3

@@ 1 Introduction
fMRI (functional MRI) has emerged as a powerful tool for measuring brain activity. This enables brain scientists to explore active brain areas during various functions and behaviors (kanwisher1997fusiform; epstein1998cortical; downing2001cortical; tang2017using; heeger2002does). However, a human can spend only limited time inside an fMRI machine. This results in fMRI-datasets too small to span the huge space of brain functionality or visual stimuli (natural images). Moreover, the variability in brain structure and function responses between different people (riddle1995individual; frost2012measuring; conroy2013inter; zhen2015quantifying) makes it difficult to combine data across individuals that have not been exposed to the same stimuli. All of these form severe limitations on current ability to analyze brain functionality.
@@
Image-to-fMRI encoding models, which predict fMRI responses to natural images, have greatly advanced the field. With the rise of deep learning, sophisticated encoding models have emerged  (yamins2014performance; eickenberg2017seeing; wen2018neural; wen2018deep; beliy2019voxels; gaziv2022self), offering novel insights into brain function (tang2024brain; henderson2023texture; gu2023human; luo2024brain; luo2023brainscuba). However, these models are primarily subject-specific and machine-specific, requiring extensive individual data (which is prohibitive) for effective training. This limits the practical use of existing brain-encoders, and prevents their ability to leverage cross-subject data. Attempts to create multi-subject encoders (e.g., (van2018modeling; khosla2020shared; wen2018transferring; gu2022personalized)) have so far been very restrictive (see Sec. 2), showing no generalization across different datasets/machines.
@@ 2 Related Work
Methods for Mapping visual stimuli to brain activity (“Image-to-fMRI Encoding”), have significantly advanced over the years, with contributions to Neuroscience. Initially, these models utilized linear regression between hand-crafted image features, to predict fMRI responses on images (kay2008identifying; naselaris2011encoding). Over time, the field has evolved to incorporate deep learning approaches both for image feature extraction and training (yamins2014performance; eickenberg2017seeing; wen2018neural; wen2018deep; beliy2019voxels; gaziv2022self; wang2023better; takagi2023high; adeli2023predicting). These models are typically subject-specific, due to substantial differences between brain responses of different people (riddle1995individual; frost2012measuring; conroy2013inter; zhen2015quantifying). This restricts the encoder to subject-specific data, and limits its generalization across subjects.
Multi-Subject Brain Models: Efficiently integrating data from multiple subjects is challenging due to anatomical and functional differences between brains. While anatomical alignment (mazziotta2001probabilistic; talairach19883; fischl2012freesurfer; dale1999cortical) aligns brains to a common anatomical space, it fails to provide accurate functional correspondences(mazziotta2001probabilistic; haxby2011common; yamada2015inter; brett2002problem; wasserman2024functional). Although we deal here with Multi-Subject Brain Encoding, we review both Multi-Subject Encoding & Decoding.
@@
In fMRI-to-Image decoding, the goal is to reconstruct an image given its corresponding fMRI brain recording. Recent advanced fMRI decoders leverage multi-subject data (e.g., ‘MindBridge’ (wang2024mindbridge), ‘MindEye2’ (scotti2024mindeye2), (gong2024mindtuner) and (han2024mindformer)), by projecting the subject-specific fMRIs of all subjects into a shared space, which is then fed into a shared image-decoding network. However, the problems of multi-subject fMRI Decoding & Encoding are fundamentally different. In multi-subject Decoding, the network inputs reside in different spaces (individual subject-specific fMRIs), whereas the network outputs reside in a shared space (the Image-Space). Thus projecting to a shared space before decoding is a valid solution. In contrast, in multi-subject Encoding, the inputs (images) reside in a shared space, whereas the outputs (individual subject-specific fMRIs) reside in completely different spaces. Thus the above approach of ‘MindBridge’&‘MindEye2’ will not apply to multi-subject Encoding, as projecting to a shared space will lose most subject-specific fMRI variability (output variability). This is resolved by our new voxel-centric model.
@@
Few attempts have been made to leverage multi-subject data for Image-to-fMRI Encoding. van2018modeling integrated data from multiple subjects, but required them to be exposed to same images, which is very restrictive. Our approach requires NO shared data between subjects. wen2018transferring and gu2022personalized finetune models of one subject to another. This allows per subject adaption, but largely limits the information sharing as they are not trained jointly. Lastly, khosla2020shared use a model with part shared and part subject-specific. However, to obtain good results they predict only a small subset of voxels which have high correlation across brains.
@@ 3.1 Overview of the Approach
The shared network components (shared across all brain-voxels of all subjects) include the feature extraction block (Fig2a), and the voxel-image cross-attention block (Fig.2c). Given any image, our encoder extracts from it various image features (from low-level to high-level) using a DINO-v2 (oquab2023dinov2) adapted model as our feature extraction block. It outputs image features from different intermediate layers of DINO, allowing each brain-voxel to attend to the appropriate semantic levels of features that align with its functionality. These image features, along with a specific voxel-embedding, are processed through the cross-attention block, to effectively integrate and predict the voxel’s activation in response to the given image. Note that this architecture is indifferent to the number of brain-voxels in each fMRI scan, hence applies to any fMRI scan.
@@ 3.2 Architecture and Training
(Fig. 2a). This block utilizes an adapted DINO model (oquab2023dinov2) to derive multi-scale image features. Features are extracted from L=5 intermediate layers of the DINO-v2 VIT-L/14 model (layers 1,6,12,18,24), where lower layers capture low-level image features and higher layers provide more semantic information. This hierarchical feature extraction is crucial, as voxels in the visual cortex correspond to a range of image attributes – from simple visual details to complex semantic content. Each layer’s features are projected to a lower-dimension C (via a linear layer), and are then concatenated along another dimension of length L (the number of layers). Since DINO operates on P image patches, the final feature output is of size L×P×C. In order to transform Dino features into features suitable for predicting brain activity, we used a LoRA inspired approach (hu2021lora), that is more suitable for data-limited settings (see A.2 for details).
@@
[…]
@@ 4 Experiments & Evaluations
We experimented with 3 prominent fMRI datasets: (i) The old “vim1” dataset (kay2008identifying; naselaris2009bayesian; kay2011fmri), which has 1750 train and 120 test grayscale images, and their corresponding 4-Tesla fMRI scans for 2 subjects. (ii) The “fMRI-on-ImageNet” scans (horikawa2017generic) comprising 1200 train and 50 test pairs of natural images from ImageNet, with 3-Tesla fMRI recordings on 5 subjects. (iii) The ”Natural Scenes Dataset” (NSD) (allen2022massive), a new 7-Tesla dataset with 8 subjects, each having around 9000 unique subject-specific images, and $\sim$1000 images shared across all subjects (which we use as our test set). This results in a total of 73,000 images, all taken from the COCO dataset (deng2009imagenet). A few example images from each of the 3 datasets are displayed in Fig. 1. A subset of visually sensitive brain voxels was either provided by the dataset or selected based on SNR (see Appendix A.1)
@@
Figure 4: The Wisdom of a Crowd of Brains. By aggregating data from multiple subjects, our Universal-Encoder improves encoding of any individual subject. We compared 3 models: (i) ”Baseline” single-subject encoder of (gaziv2022self), (ii) ”Universal Encoder - single subject” – our architecture trained on each subject separately, (iii) ”Universal Encoder - multiple subjects” – our model trained on data from 8 subjects. (a) Pearson Correlation (per voxel) between predicted & ground-truth fMRI (Median value, 75th & 25th percentiles). (b) Retrieval Accuracy (Top-1 & Top-5) of the GT image per “Query” fMRI.
@@ 4.1 The Wisdom of a “Crowd of Brains”
We first demonstrate the Universal-Encoder’s ability to exploit data from multiple subjects, without any shared-data. For this we use the 7-Tesla NSD dataset. We train on union of all 8 subject-specific images ($\sim$72,000 non-shared data), and test on the $\sim$1,000 shared images (the images that all 8 subjects saw). We compared 3 models in our evaluation: (i) As a baseline, we used the image-to-fMRI encoder of (gaziv2022self), trained separately for each subject on their subject-specific training-set (“Baseline encoder”). (ii) Our Universal-Encoder trained on each subject separately (“Universal Encoder - single subject”), and (iii) Our Universal-Encoder trained on all 8 subjects jointly, using their combined training-sets, and tested on each subject individually (“Universal Encoder - multiple subjects”). A comparison to two other prominent encoder models (wang2023better; takagi2023high) (which have lower performance), is further provided in the appendix Fig. S8.
Fig. 4a shows the median Pearson correlation value (along with the $25$th & $75$th percentiles), computed between the predicted activations and the ground-truth activations for all fMRI voxels. These are computed per subject, for each of the 3 models. Our Universal-Encoder, even when trained on a single subject, performs consistently better than the Baseline-encoder. Moreover, when our Universal-Encoder is trained jointly on the training sets of all subjects, it consistently outperforms all subject-specific models. It obtains notable improvements for both the “best” subjects (Subject1: $\sim$7% improvement) and the “worst” subject (Subject8: $\sim$15% improvement). Statistical significance of the performance gap is assessed in Appendix C, confirming significant superiority of the Universal-Encoder. Per-region results are further presented in Fig. S10.
@@
Fig. 4b presents quantitative Retrieval results evaluated per subject for all 3 models. It shows the percent of times the correct image (corresponding to the ”Query” fMRI) was ranked 11st (Top-1) and among the Top-5 retrieved images. The results indicate superior performance of the multi-subject Universal-Encoder compared to the 2 other models in both Top-1 & Top-5. Statistical significance of the Universal Encoder’s improvement over other models is verified in Appendix C & Table.S2. Our experiments demonstrate the model’s ability to effectively aggregate data from multiple subjects (who viewed different images), while enhancing the performance of each subject individually.
@@ 4.2 The Wisdom of a “Crowd of Datasets”
Fig. 5a compares 3 models for the “fMRI-on-ImageNet” dataset: (i) Single-subject encoder of (gaziv2022self) (“Baseline encoder”), (ii) Our Universal-Encoder trained on all subjects in “fMRI-on-ImageNet” (“Universal Encoder - same dataset”), and (iii) Universal-Encoder trained on subjects from both “fMRI-on-ImageNet” and NSD (“Universal Encoder - multi datasets”). Our multi-subject same-dataset encoder (green) outperforms the single-subject baseline model (blue). Adding data from NSD yields even further improvement (red). Median correlation, 75th & 25th percentiles are shown. Fig. 5b shows results for VIM1 dataset. Here too, adding training data from NSD significantly enhances encoding performance on VIM1.
@@ 4.3 Transfer-Learning to New Subjects/Datasets
The transferred Universal-Encoder significantly outperforms any single-subject model on all datasets. Transfer-Learning to Subject1 within NSD (Fig. 6a) obtains $\geq$77% improvement for any number of training examples. Moreover, as few as 100 subject-specific training examples suffice to obtain better results than a subject-specific model trained on the entire train-set (9000 examples). A similar gap in performance is achieved for Subject8.
@@
Fig. 6b (”fMRI-on-ImageNet”) & Fig. 6c (VIM1) demonstrate transfer learning from a new 7-Tesla dataset to older lower-resolution 3T or 4T datasets. These were scanned on very different types of images (e.g., VIM1 was scanned on B/W images with a circular black mask), and have much smaller train-sets. Transfer-Learning shows a significant improvement for any number of training data, with an improvement $\geq$84% for ‘fMRI-on-ImageNet’ & $\sim$45% for VIM1, using the entire (small) subject-specific train-set (significance testing over model initialization in Table. S3).
@@ Appendix A Further Technical Details
The datasets utilized in our study comprise BOLD fMRI responses to various natural images, recorded over multiple scanning sessions. Peak BOLD responses corresponding to each stimulus were estimated, resulting in a scalar value for each voxel for each image. Each dataset underwent unique pre-processing procedures, detailed in their respective publications (kay2008identifying; naselaris2009bayesian; Horikawa2017GenericFeatures; allen2022massive).
@@
For the NSD dataset (allen2022massive), we used a post-processed dataset with voxel selection from (gifford2023algonauts). They chose vision-related areas, resulting in a total of approximately 40,000 voxels, and provided voxels after per voxel normalization.
@@
For the fMRI-on-ImageNet dataset (horikawa2017generic), a relevant set of around 5,000 voxels was already provided. We implemented Z-scoring normalization ourselves for this dataset.
@@
For the VIM1 dataset (kay2008identifying; naselaris2009bayesian), we selected the best 7,000 voxels according to the highest Signal-to-Noise Ratio (SNR). SNR is calculated as the ratio of the variance of averaged (repeated) measurements for different stimuli to the average variance of measurements for the same stimuli. This approach ensures the selection of voxels most representative of neural activity in response to diverse visual stimuli. Z-scoring normalization was also implemented for this dataset.